# Causal Attribution via Activation Patching

### 13 Mar 2026 v1

@@ Abstract
Attribution methods for Vision Transformers (ViTs) aim to identify image regions that influence model predictions, but producing faithful and well-localized attributions remains challenging. Existing gradient-based and perturbation-based techniques often fail to isolate the causal contribution of internal representations associated with individual image patches. The key challenge is that class-relevant evidence is formed through interactions between patch tokens across layers, and input-level perturbations can be poor proxies for patch importance, since they may fail to reconstruct the internal evidence actually used by the model. We propose Causal Attribution via Activation Patching (CAAP), which estimates the contribution of individual image patches to the ViT’s prediction by directly intervening on internal activations rather than using learned masks or synthetic perturbation patterns. For each patch, CAAP inserts the corresponding source-image activations into a neutral target context over an intermediate range of layers and uses the resulting target-class score as the attribution signal. The resulting attribution map reflects the causal effect of patch-associated internal representations on the model’s prediction. The causal intervention serves as a principled measure of patch influence by capturing class-relevant evidence after initial representation formation, while avoiding late-layer global mixing that can reduce spatial specificity. Across multiple ViT backbones and standard metrics, CAAP significantly outperforms existing methods and produces more faithful and localized attributions.
@@ 1 Introduction
Vision Transformers (ViTs) have become a dominant architecture for image recognition and representation learning, including supervised and foundation models such as ViT [34], CLIP [27], DINOv2 [22] and DeiT3 [37]. Their patch-based tokenization and global self-attention naturally motivate attribution methods that operate at the patch or token level rather than on individual pixels [9]. Input attribution methods aim to quantify how specific input features contribute to a model’s prediction and are commonly used to analyze model behavior, identify reliance on spurious cues, and compare learned representations [29]. Existing attribution approaches for vision models include gradient-based methods [31, 35, 30], perturbation-based methods [25, 40, 8], and combined methods that use gradients and attention [7, 26]. Despite advances, there is still room for improvement in producing faithful and well-localized attributions for ViTs.
@@
Perturbation-based methods, including masking approaches in input or token space [25, 40, 8], estimate importance by measuring prediction changes under interventions and therefore provide a causal notion of attribution. However, they have important practical limitations. Deletion-based evaluation does not account for redundancy across patches: when multiple patches encode similar class evidence, removing one patch may cause little change in the output, leading to an underestimation of its importance. Insertion-based approaches face a different limitation. Inserting a single patch into an otherwise blank or blurred input fails to reproduce the functional role that the patch plays within the full image context. Moreover, input-space perturbations can introduce distribution shift, and existing methods often rely on input masks [10, 40] or optimization-based [38] surrogates that may deviate from the model’s native internal decision variables. These limitations motivate causal attribution methods that operate directly on internal representations rather than on raw input [6, 42].
@@
In this work, we propose Causal Attribution via Activation Patching (CAAP), a novel attribution method motivated by mechanistic interpretability. CAAP measures attribution by intervening on hidden activations of a trained ViT. For each patch, we insert the corresponding source-image activations into a neutral blank context and use the resulting target-class score as the attribution signal. Crucially, we apply activation patching over an intermediate layer range: late enough that patch representations contain enriched class-relevant object information, but early enough to avoid late-layer contextual mixing that can reduce spatial specificity. This design allows CAAP to better isolate the causal contribution of class-relevant patch representations to the final prediction.
@@
Causal intervention offers a practical advantage for CAAP over existing attribution techniques. It assigns attribution scores directly based on information already encoded in the internal representations of tokens, rather than approximating relevance through gradients, attention weights, or learned saliency masks. The procedure can be efficiently parallelized, enabling practical patch-level attribution for large ViT backbones. Empirical results across multiple Vision Transformer backbones show that our method significantly outperforms existing approaches, achieving state-of-the-art results on deletion, insertion, and foreground and background localization metrics. Quantitative and qualitative results demonstrate that this method provides a reliable and effective mechanism for spatial attribution in Vision Transformers, introducing a new methodological paradigm for attribution, based on causal interventions.
@@ 2.1 Problem Definition
We consider an image classification model $f$ that maps an input image $x\in\mathbb{R}^{C\times H\times W}$ to a vector of class scores $f(x)\in\mathbb{R}^{K}$, where $K$ is the number of classes. For a target class $y$, we denote the corresponding prediction score by $p(y\mid x)=f(x)_{y}$. The image $x$ is partitioned into a set of non-overlapping visual patches $\mathcal{P}=\{p_{1},\dots,p_{N}\}$, where each patch corresponds to a spatial region in the input. The goal of visual attribution is to assign a scalar attribution score $A(p_{i})\in\mathbb{R}$ to each patch $p_{i}$, measuring the contribution of that patch to the prediction $f(x)_{y}$. An attribution method produces an attribution map $A\in\mathbb{R}^{N}$ such that higher values indicate greater relevance to the target prediction.
@@ 2.2 Activation Patching
while all remaining activations are propagated according to the model’s standard forward computation. The model output $f(c^{*})$ reflects the effect of intervening on the selected internal activations of $c$, enabling causal analysis of their influence on the final prediction.
@@ 3 Method
We propose CAAP for ViT models, a patch-level attribution method that assigns relevance scores by intervening on the internal activations of a trained classifier. Rather than measuring local sensitivity at the input level, CAAP quantifies the causal effect of internal representations associated with a visual patch on the final prediction. This addresses a core limitation of input-level insertion and deletion methods. Deletion can underestimate importance when class-relevant information is redundant and spread across multiple patches, while insertion into a blank image ignores token interactions across transformer layers.
@@
This score measures how predictive the contextualized token embedding associated with patch $p_{i}$ is for the target class, by inserting it into a neutral context and observing its effect on the output. Repeating this procedure for all patches yields a patch-level attribution map over the input image. The overall procedure is summarized in Algorithm 1 and shown in Fig. 1.
@@ 3.2 Choice of Intervened Layers $\mathcal{L}$
Vision Transformers build patch representations through repeated self-attention, causing each patch token to encode information aggregated from other patches. This perspective suggests distinguishing two representation regimes in ViTs [23]. First, patch tokens undergo a grouping process in which class-consistent and object-related features become organized across patches. Second, tokens increasingly encode broader context through interactions that may involve tokens from different objects or background regions. For most pretrained ViTs, intra-object grouping occurs in the early layers, while inter-object or background contextualization emerges in the later layers [23].
@@ 3.3 Efficient Implementation
In transformer architectures, the hidden states at each layer are linearly projected to queries, keys, and values. To avoid recomputation of these vectors from hidden states during activation patching, we reuse the cached key and value projections associated with the intervened hidden states rather than recomputing them from the residual streams. Only the query vectors of the additional $\mathrm{CLS}$ tokens are computed dynamically. We next propose two practical ideas to further accelerate the intervention process.
3.3.1 Parallel computation over patches
@@
The patch-wise formulation requires one forward pass per patch, which scales linearly with the number of tokens in the image. Since the intervention is defined on internal activations, we can compute all patch attribution scores in parallel within a single forward pass while preserving the same causal intervention. This construction exploits the structure of activation patching and is not applicable to input-space masking methods, which require independent forward passes for each modified input.
@@
For the blank image $x_{0}$, we perform a single forward pass once and cache its activations $\{h^{l}(x_{0})\}_{l\in\mathcal{L}}$, which are reused for all image samples. Then, for each source image $x$, we run a standard forward pass and cache the activations $\{h^{l}(x)\}_{l\in\mathcal{L}}$. We then build a patched computation that contains both activation sets. For each patch $p_{i}$, we introduce a dedicated classification token $\mathrm{CLS}_{i}$ with the same positional encoding as the original classification token. The attention mask is defined such that $\mathrm{CLS}_{i}$ attends to source activations at indices $\mathcal{S}(p_{i})$ and to target activations at indices $\neg\mathcal{S}(p_{i})$. No attention is allowed between different classification tokens, and only the $\mathrm{CLS}$ tokens are propagated through the remaining layers.
The class score produced by $\mathrm{CLS}_{i}$ for class $y$ is equal to the attribution score in Eq. 3 obtained by patch-wise activation replacement for $p_{i}$. Hence, this construction yields the same attribution map while computing all patch scores in parallel. Therefore, the total computational cost per dataset sample consists of one forward pass for the source image, followed by a single forward pass with multiple $\mathrm{CLS}$ tokens. Since the additional computation arises only from extra query tokens with shared key–value attention, the cost is bounded by a small constant factor relative to a standard forward pass.
3.3.2 Precomputation of blank-context attention
Even in the parallel construction, a large fraction of the computation arises from attention between each classification token and patch tokens originating from the blank image $x_{0}$. Let $K_{0}^{l}$ and $V_{0}^{l}$ denote the key and value matrices at layer $l$ computed from $x_{0}$. These quantities are fixed and do not depend on the source image $x$ or on the selected patch $p_{i}$. In the patched context $c^{*}$, the update of each classification token $\mathrm{CLS}_{i}$ follows the same self-attention operation defined in Eqs. 4 and 5, with the query corresponding to $\mathrm{CLS}_{i}$ rather than a patch token. In this setting, keys and values are drawn either from source-patched tokens $j\in\mathcal{S}(p_{i})$ or from blank-context tokens $j\notin\mathcal{S}(p_{i})$.
Empirically, we observe that for blank patches the dot products $\langle q_{\mathrm{CLS}}^{c^{\prime}},k_{0,j}^{l}\rangle$ and $\langle q_{\mathrm{CLS}}^{c^{*}(p_{i})},k_{0,j}^{l}\rangle$ are nearly identical across images and interventions. As a result, the attention distribution over blank tokens is effectively invariant to the inserted source activations. We therefore precompute, for each layer and head, the aggregated contribution of blank patches to the $\mathrm{CLS}$ update using the corresponding statistics from the blank context $c^{\prime}$ (see Appendix B.1 for more details). During attribution, we reuse these fixed quantities and compute only the terms involving source-patched tokens dynamically. This avoids repeated evaluation of dot products and attention weights for blank patches across images. In practice, this approximation does not reduce faithfulness or localization performance, while decreasing floating-point operations by approximately an order of magnitude.
@@ 4 Experiments
In this section we present qualitative and quantitative evaluations of our proposed ViT attribution method. We begin by defining the evaluation metrics used throughout the experiments, followed by a detailed description of the ViT models, datasets, and baseline methods. We then present the quantitative results and ablation studies analyzing the contribution of each component of our method.
@@ 4.1 Evaluation Metrics
We evaluate attribution maps using six metrics that assess two complementary aspects of quality: Faithfulness and Localization. Deletion AUC and Insertion AUC [25] are faithfulness measures; they test whether pixels ranked as important by the attribution are indeed the ones the model relies on. To obtain a single faithfulness metric, we also report Insertion-Deletion, defined as the difference between Insertion and Deletion AUC. AUPR1, AUPR0 [19], and Pointing-Game (PG) [3] are localization measures; they quantify how well attribution mass concentrates on annotated foreground regions and avoids background regions.
@@
Following the evaluation protocol of [25], given an attribution map $A$ for an image $x$ and corresponding class $y$, we sort pixels in descending order of $A$ and perturb $x$ accordingly. For Deletion, we progressively replace the top-ranked pixels with a constant reference value and track the target class score $f(\cdot)_{y}$; the area under this curve is Deletion AUC. For Insertion, we start from a blurred version of $x$ and progressively reveal (unblur) the top-ranked pixels, again tracking $f(\cdot)_{y}$; the area under this curve is Insertion AUC. More faithful attributions yield lower Deletion AUC and higher Insertion AUC.
@@ 4.2 Experiment Settings
We conduct our main experiments on the ImageNet-1K (ILSVRC2012) validation set [28]. We evaluate four ViT-L models on ImageNet-1K: CLIP-L/14 [27], DINOv2-L/14 [22], ViT-L/16 [34], and DeiT3-L/16 [37]. For localization evaluation on ImageNet-1K, we use ImageNet-S [13] annotations and report AUPR1, AUPR0, and PG. Results for additional architectures are provided in Appendix A. To further validate our method, we additionally evaluate on ImageNet-Hard [36], Oxford-IIIT-Pet [24], and Food-101 [4]. For these datasets, we use a ViT-B/16 classifier initialized from pretrained weights. Results are reported on a random sample of 500 validation images per dataset.
@@
We compare our approach against a strong set of widely-used attribution baselines drawn from various attribution categories, including Grad-CAM [30], Attention Rollout [1], Transformer Attribution (T-Attr) [7], ViT-CX [40], TAM [41], IxG+ [21], Libra IxG+ [20], MDA [38], and MUTEX [18]. We evaluate all methods using the same protocol and metrics as our method.
4.3 Results

### 17 May 2026 v2

@@ Abstract
Attribution methods for Vision Transformers (ViTs) aim to identify image regions that influence model predictions, but producing faithful and well-localized attributions remains challenging. Existing attribution methods face several limitations, with gradient-based, relevance-propagation, and attention-based methods relying on local approximations, while perturbation or optimization-based methods intervene on inputs, tokens, or surrogates rather than internal patch representations. The key challenge is that class-relevant evidence is formed through interactions between patch tokens across layers; methods that operate only on input changes, attention weights, or backward relevance signals may therefore provide indirect proxies for patch importance rather than directly testing the predictive effect of contextualized patch representations. We propose Causal Attribution via Activation Patching (CAAP), which estimates the contribution of individual image patches to the ViT’s prediction by directly intervening on internal activations rather than using learned masks or synthetic perturbation patterns. For each patch, CAAP inserts the corresponding source-image activations into a neutral target context over an intermediate range of layers and uses the resulting target-class score as the attribution signal. The resulting attribution map reflects the causal contribution of patch-associated internal representations on the model’s prediction. The causal intervention serves as a principled measure of patch influence by capturing semantic evidence after initial representation formation, while avoiding late-layer global mixing that can reduce spatial specificity. Across multiple ViT backbones and standard metrics, CAAP consistently outperforms existing methods in various settings and produces more faithful and localized attributions.
@@ 1 Introduction
Vision Transformers (ViTs) have become a dominant architecture for image recognition and representation learning, including supervised and foundation models such as ViT [35], CLIP [28], DINOv2 [23] and DeiT3 [38]. Their patch-based tokenization and global self-attention naturally motivate attribution methods that operate at the patch or token level rather than on individual pixels [9]. Input attribution methods aim to quantify how specific input features contribute to a model’s prediction and are commonly used to analyze model behavior, identify reliance on spurious cues, and compare learned representations [30]. Existing attribution approaches for vision models include gradient-based and relevance-propagation methods [32, 36, 31, 7, 22, 21], attention-flow and token-mixing methods [1, 42, 27, 19], and perturbation or optimization-based methods that modify inputs, tokens, attention mechanisms, or patch embeddings [26, 12, 11, 10, 8, 41, 39]. Despite this broad progress, producing attribution maps that are both faithful to ViT predictions and spatially localized remains challenging.
@@
A central difficulty in ViT attribution is that patch relevance is not determined only by the raw visual content of a patch, but by contextualized representations formed through self-attention across layers. This creates limitations for several attribution families. Gradient-based and relevance-propagation methods provide efficient class-specific signals, but they measure local sensitivity or propagate relevance through the computation graph rather than directly testing whether a contextualized patch representation is sufficient to support the prediction [32, 36, 31, 7, 21]. Attention-based methods expose token-mixing patterns, but attention weights or attention flow are indirect proxies for feature contribution, especially after information from different tokens has been repeatedly mixed across layers [1, 42, 27].
@@
Perturbation and masking methods measure causal output changes under interventions, but input-space perturbations can introduce distribution shift [26, 12, 11, 10]. Input-based deletion does not account for redundancy across patches: when multiple patches encode similar class evidence, removing each patch from a group of redundant patches may cause little change in the output, leading to underestimation of the group’s importance. Insertion into a blank or blurred image overlooks interactions among patches and may fail to reproduce the functional role a patch plays in the full image context. More ViT-specific perturbation or optimization methods operate over tokens, attention mechanisms, patch embeddings, or attribution metrics [8, 41, 39], but they still do not directly evaluate the class evidence carried by the original internal patch representation within a controlled target context.
@@
In this work, we propose Causal Attribution via Activation Patching (CAAP), a novel attribution method motivated by mechanistic interpretability. Rather than attributing through backward sensitivity, attention propagation, input masking, or optimized surrogate maps, CAAP directly intervenes on the hidden activations of a trained ViT. For each patch, we insert the corresponding source-image activations into a neutral blank context and use the resulting target-class score as the attribution signal [43]. Crucially, we apply activation patching up to the intermediate layers: late enough that patch representations contain enriched class-relevant object information, but early enough to avoid late-layer contextual mixing which reduces spatial specificity. This design allows CAAP to better isolate the causal contribution [14, 6] of class-relevant patch representations to the final prediction.
@@
Causal intervention offers a practical advantage for CAAP over existing attribution techniques. It derives patch scores directly from information already encoded in contextualized token representations, rather than relying on indirect evidence such as attention weights, approximate gradient signals, input-space masks, or metric-optimized surrogate maps. The procedure can be efficiently parallelized, enabling practical patch-level attribution for large ViT backbones. Empirical results across multiple Vision Transformer backbones show that our method improves over existing approaches on most evaluated settings, achieving state-of-the-art results on deletion, insertion, and foreground and background localization metrics. Quantitative and qualitative results demonstrate that this method provides a reliable and effective approach to spatial attribution in ViTs, offering a new perspective based on mechanistic interpretability tools.
@@ 2.1 Problem Definition
We consider an image classification model $f$ that maps an image $x\in\mathbb{R}^{C\times H\times W}$ to the class scores $f(x)\in\mathbb{R}^{K}$, where $K$ is the number of classes. For a target class $y$, we denote the corresponding prediction score by $p(y\mid x)=f(x)_{y}$. The image $x$ is partitioned into a set of non-overlapping visual patches $\mathcal{P}=\{p_{1},\dots,p_{N}\}$, where each patch corresponds to a spatial region in the input. The goal of visual attribution is to assign a scalar attribution score $A(p_{i})\in\mathbb{R}$ to each patch $p_{i}$, measuring the contribution of that patch to the prediction $f(x)_{y}$. An attribution method produces an attribution map $A\in\mathbb{R}^{N}$ such that higher values indicate greater relevance to the target prediction.
@@ 2.2 Activation Patching
while all remaining activations are propagated according to the model’s standard forward computation. The model output $f(c^{*})$ reflects the effect of intervening on the selected internal activations of $c$, enabling causal analysis of their influence on the final prediction under activation interventions [14].
@@ 3 Method
We propose CAAP for ViT models, a patch-level attribution method that assigns relevance scores by intervening on the internal activations of a trained classifier. Rather than measuring local sensitivity at the input level, CAAP quantifies the causal effect of internal representations associated with a visual patch on the final prediction. This addresses a core limitation of input-level insertion and deletion methods. Deletion-based methods can underestimate importance when class-relevant information is redundant and spread across multiple patches, and insertion into a blank image produces an out-of-distribution (OOD) input and ignores token interactions across transformer layers.
@@
This score measures how predictive the contextualized token embedding associated with patch $p_{i}$ is for the target class, by inserting it into a neutral context and observing its effect on the output. Repeating this procedure for all patches yields a patch-level attribution map over the input image.
An important property of activation patching, compared to input-based perturbation methods, is that the patched activations are taken directly from the original context, so they already encode interactions between patches and avoid introducing OOD inputs in the source context. The overall procedure of the proposed method is summarized in Algorithm 1 and shown in Fig. 1.
@@ 3.2 Choice of Intervened Layers $\mathcal{L}$
ViTs build patch representations through repeated self-attention, causing each patch token to encode information aggregated from other patches. This perspective suggests distinguishing two representation regimes in ViTs [24]. First, patch tokens undergo a grouping process in which class-consistent and object-related features become organized across patches. Second, tokens increasingly encode broader context through interactions that may involve tokens from different objects or background regions. For most pretrained ViTs, intra-object grouping occurs in the early layers, while inter-object or background contextualization emerges in the later layers [24].
@@ 3.3 Efficient Implementation
In transformers, hidden states at each layer are linearly projected to queries, keys, and values. To avoid recomputing these vectors from hidden states during activation patching, we reuse cached key and value projections associated with the intervened hidden states, rather than recomputing them from residual streams. Only the query vectors of the additional $\mathrm{CLS}$ tokens are computed dynamically in the patched context. We next propose two practical techniques to further accelerate the intervention process: parallel computation of $\mathrm{CLS}$ tokens and precomputation of blank-context attention.
@@
The patch-wise formulation requires one forward pass per patch, which scales linearly with the number of tokens in the image. Since the intervention operates on internal activations, we can compute all patch attribution scores in parallel within a single forward pass while preserving the same causal intervention. This construction exploits the structure of activation patching and is not applicable to input-space masking methods, which require independent forward passes for each modified input.
@@
For the blank image $x_{0}$, we perform one forward pass and cache its activations $\{h^{l}(x_{0})\}_{l\in\mathcal{L}}$, reused across samples. For each source image $x$, we run a standard forward pass and cache $\{h^{l}(x)\}_{l\in\mathcal{L}}$. We then build a patched computation containing both activation sets. For each patch $p_{i}$, we introduce a dedicated classification token $\mathrm{CLS}_{i}$ with the same positional encoding as the original $\mathrm{CLS}$. The attention mask ensures that $\mathrm{CLS}_{i}$ attends to source activations at $\mathcal{S}(p_{i})$ and to target activations at $\neg\mathcal{S}(p_{i})$. No attention is allowed between different classification tokens, and only the $\mathrm{CLS}$ tokens are propagated through the remaining layers. The resulting class score from $\mathrm{CLS}_{i}$ for class $y$ equals the attribution score in Eq. 3 from patch-wise activation replacement for $p_{i}$.
This parallel construction of $\mathrm{CLS}$ tokens yields the same attribution map (no approximation) while computing all patch scores in parallel. The total cost per sample is one forward pass for the source image and one forward pass with multiple $\mathrm{CLS}$ tokens. In addition, we introduce an extra efficiency method in Appendix C, based on the fact that a large fraction of the computation comes from attention between each $\mathrm{CLS}$ token and blank-context tokens. We exploit the near-invariance of these interactions to precompute their contribution and reuse it during attribution. This approximation reduces computation without notable effect on performance (details and results are in Appendix C).
@@ 4 Experiments
In this section, we present qualitative and quantitative evaluations of CAAP. We begin by defining the evaluation metrics used throughout the experiments, followed by a detailed description of the ViT models, datasets, and baseline methods. We then present qualitative and quantitative results, along with ablation studies analyzing the contribution of each individual component of CAAP.
@@ 4.1 Evaluation Metrics
We evaluate attribution maps using six metrics that assess two complementary aspects of quality: faithfulness and localization. Deletion AUC and Insertion AUC [26] are faithfulness measures; they test whether pixels ranked as important by the attribution are indeed the ones the model relies on. To obtain a single faithfulness metric, we also report Insertion-Deletion, defined as the difference between Insertion AUC and Deletion AUC. AUPR1, AUPR0 [20], and Pointing-Game (PG) [3] are localization measures; they quantify how well attribution mass concentrates on annotated foreground regions and avoids background regions. See Appendix B.2 for additional metrics that are less common in the literature.
@@
Following the evaluation protocol of [26], given an attribution map $A$ for an image $x$ and corresponding class $y$, we sort pixels (patches) in descending attribution order $A$ and perturb $x$ accordingly. For Deletion, we progressively replace the top-ranked pixels with a constant reference value and track the target class score $f(\cdot)_{y}$; the area under this curve is Deletion AUC. For Insertion, we start from a blurred version of $x$ and progressively reveal (unblur) the top-ranked pixels, again tracking $f(\cdot)_{y}$; the area under this curve is Insertion AUC. More faithful attributions yield lower Deletion AUC, higher Insertion AUC and higher Insertion-Deletion.
@@ 4.2 Experiment Settings
We conduct our main experiments on the ImageNet-1K (ILSVRC2012) validation set [29]. We evaluate four ViT-L models on ImageNet-1K: CLIP-L/14 [28], DINOv2-L/14 [23], ViT-L/16 [35], and DeiT3-L/16 [38]. For localization evaluation on ImageNet-1K, we use ImageNet-S [13] annotations and report AUPR1, AUPR0, and PG. Results for additional architectures are provided in Appendix B. To further validate CAAP, we additionally evaluate on ImageNet-Hard [37], Oxford-IIIT-Pet [25], and Food-101 [4]. For these datasets, we use a ViT-B/16 classifier initialized from pretrained weights.
@@
We compare CAAP against a strong set of widely-used attribution baselines drawn from various attribution categories: gradient-based methods, including Grad-CAM [31], IxG+ [22], and Libra IxG+ [21]; attention-based methods, including Attention Rollout [1], Transformer Attribution (T-Attr) [7], TAM [42], and MUTEX [19]; and perturbation-based methods, including RISE [26], ViT-CX [41], and MDA [39]. We evaluate all methods using the same protocol and metrics.
4.3 Quantitative Results
Faithfulness.
Table 1 reports faithfulness metrics across four ViT models. Table 3 further presents the same faithfulness metrics on ImageNet-Hard [37] and Oxford-IIIT-Pet [25]. CAAP consistently achieves the lowest Deletion AUC and the highest or competitive Insertion AUC on all backbones. In particular, CAAP improves the Insertion-Deletion with a notable margin compared to prior methods, indicating that the regions ranked as important by our attribution maps more accurately reflect the model’s true decision-making process.